OpenStudy (anonymous):

if q/r=3 and r/t=5 and t/u=7. what is the value of qu/t^2?

OpenStudy (anonymous):

\[\frac{q}{r}=3 , \frac{r}{t}=5,\frac{t}{u}=7\] \[\frac{u}{t}=\frac{1}{7}\] so how could you combine all of this to get \[\frac{qu}{t^2}\]

OpenStudy (anonymous):

I'm still a bit confused. @completeidiot

OpenStudy (anonymous):

ok what happens if you multiply \[\frac{q}{r}*\frac{r}{t}\]?

OpenStudy (anonymous):

qt=r^2

OpenStudy (anonymous):

there was no equal sign

OpenStudy (anonymous):

Oh right, careless mistake. qr/rt

OpenStudy (anonymous):

and the r cancels out right?

OpenStudy (anonymous):

yes @completeidiot

OpenStudy (anonymous):

so \[\frac{q}{t}=?\] if \[\frac{q}{r}=3, \frac{r}{t} = 5\]

OpenStudy (anonymous):

Do I have to plug in numbers in order to find out?

OpenStudy (anonymous):

no if we multiplied q/r by r/t the answer is the same as multiplying 3 and 5

OpenStudy (anonymous):

so q/t =?

OpenStudy (anonymous):

15..? Sorry, it's still a bit unclear to me

OpenStudy (anonymous):

3/5?

OpenStudy (anonymous):

right so now we know \[\frac{q}{t}=15\]

OpenStudy (anonymous):

\[\frac{q}{r}=3\] \[\frac{q}{r} *\frac{r}{t}=3*\frac{r}{t}\] \[\frac{q}{t}=3*5\]

OpenStudy (anonymous):

now that we have \[\frac{q}{t}\] what can we multiply to this in order to get \[\frac{qu}{t^2}\]?

OpenStudy (anonymous):

t/u

OpenStudy (anonymous):

note \[\frac{q}{t}*\frac{u}{t}=\frac{qu}{t^2}\]

OpenStudy (anonymous):

if we multiplied t/u to q/t, the t would cancel out and we would get q/u

OpenStudy (anonymous):

Oh alright, but how do you find the value?

OpenStudy (anonymous):

well we know \[\frac{t}{u}=7\] and we want \[\frac{u}{t}\] so we find the reciprocal of t/u \[\frac{1}{\frac{t}{u}}=\frac{u}{t}= \frac{1}{7}\]

OpenStudy (anonymous):

since we have q/t and u/t in order to get qu/t^2, we need to multiply them together so you just multiply the values of q/t and u/t together to get your answer